40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Mar 24 19:42:12 2014, Charles, Frogs, VAC, Power Failure 
    Reply  Mon Mar 24 21:36:28 2014, Koji, Update, General, Power Failure 
       Reply  Mon Mar 24 22:13:37 2014, steve, Update, General, Power Failure poweroutage.png
          Reply  Tue Mar 25 14:52:57 2014, steve, Update, VAC, vacuum is recovered 17hrsNoPumping.png
             Reply  Fri Mar 28 16:26:20 2014, steve, Update, VAC, RGA scan after power failure afterPowerFailure.png
          Reply  Tue Mar 25 16:11:24 2014, Koji, Update, General, Power Failure 
             Reply  Wed Mar 26 11:30:07 2014, Koji, Update, General, Power Failure 
                Reply  Wed Mar 26 22:22:43 2014, Manasa, Update, General, Recovery 
                   Reply  Thu Mar 27 10:34:39 2014, ericq, Update, General, Recovery 
Message ID: 9752     Entry time: Wed Mar 26 11:30:07 2014     In reply to: 9750     Reply to this: 9755
Author: Koji 
Type: Update 
Category: General 
Subject: Power Failure 

Recovery work: now arms are locking as usual

- FB is failing very frequently. Everytime I see red signals in the CDS summary, I have to run "sudo ntpdate -b -s -u pool.ntp.org"

- PMC was aligned

- The main Marconi returned to initial state. Changed the frequency and amplitude to the nominal value labeled on the unit

- The SHG oven temp controllers were disabled. I visited all three units and pushed "enable" buttons.

- Y arm was immediately locked. It was aligned using ASS.

- X arm did not show any flash. I found that the scx model was not successfully burtrestored yesterday.
  The setting was restored using Mar 22 snapshot.

- After a little tweak of the ETMX alignment, a decent flash was achieved. But still it could not be locked.

- Run s/LSC/LSCoffset.py. This immediately made the X arm locked.

- Checked the green alignment. The X arm green is beating with the PSL at  ~100MHz but is misaligned beyond the PZT range.
  The Y arm green is locked on TEM00 and is beating with the PSL at ~100MHz.

ELOG V3.1.3-